joshua is older than astrid. the difference of their age is 12 and the sum of their ages is 50. find the age of each.
Answer by macston(5194)

J=Joshua's age; A=Astrid's age
J-A=12
J=12+A
J+A=50 Substitute for J.
12+A+A=50 Subtract 12 from each side.
2A=38 Divide each side by 2.
A+19 ANSWER 1: Astrid is 19 years old.
J=12+A=12+19=31 ANSWER 2: Joshua is 31 years old.
CHECK:
J+A=50
31+19=50
50=50